dmcnair11:

Talia is on a road trip with some friends. In the first 2 hours, they travel 100 miles. Then they hit traffic and go only 30 miles in the next hour. The last hour of their trip, they drive 75 miles.

Vocaloid:

is this the full question? it seems like something might be missing, it's not asking anything

dmcnair11:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Vocaloid is this the full question? it seems like something might be missing, it's not asking anything \(\color{#0cbb34}{\text{End of Quote}}\) " Calculate the average speed of Talia’s car during the trip. Give your answer to the nearest whole number. ______ mph "

Vocaloid:

average speed = total distance / total time so we have total distance = 100 + 30 + 75 miles total time = 2 + 1 + 1 hours, calculate avg. speed

dmcnair11:

.........................

Vocaloid:

if something I said was unclear, you are welcome to ask questions ~